OpenStudy (anonymous):

Solve :What value of x makes the equation true? Answer(S) Two-eights Equals Two Over X 96 48 18 14 Thankssssss (;

jimthompson5910 (jim_thompson5910):

2/8 = 2/x Since the numerators are the same, and the fractions are the same, the denominators MUST be the same. So 8 = x or x = 8

OpenStudy (paxpolaris):

none of them ??? 2/8 = 2/8 two eighths = 2 over 8

jimthompson5910 (jim_thompson5910):

Alternatively, you can solve like this 2/8 = 2/x 2x = 8*2 2x = 16 x = 16/2 x = 8

OpenStudy (anonymous):

those are the answers the test gave me

OpenStudy (paxpolaris):

maybe you copied the sentence .... Two-eights Equals Two Over X ... wrong

OpenStudy (anonymous):

that wasw part of the sentce my bad

OpenStudy (paxpolaris):

what is the equation in the question ???

OpenStudy (anonymous):

What value of x makes the equation true? Two-Eights equals two over x <-- Thats the questions and I gave you the answers lol (: